OpenStudy (anonymous):

plz help a two digit num is such that the product of the digit is 12. when 36 is added to the number the digits interchange thier places. find quad eq?

OpenStudy (lgbasallote):

here's a hint.. the factors of 12 (that are consisted of 1 digit times 1 digit) are 3 x 4 4 x 3 6 x 2 2 x 6 does that help?

OpenStudy (anonymous):

find the quad equation

OpenStudy (lgbasallote):

huh? aren't you asking for the number?

OpenStudy (lgbasallote):

this question is vague

OpenStudy (lgbasallote):

very vague

OpenStudy (anonymous):

the number is 26

OpenStudy (anonymous):

let the digit in the tenth place be x then the digit in the units place is 12/x hence orig number =10x+12/x now after the digits interchange is 10*12/x + x by the problem (10x+12/x)+36=120/x +x .... and so on

OpenStudy (anonymous):

y u hve tkn 12/x . y nt 12x

OpenStudy (anonymous):

@matricked I don't think you are correct

OpenStudy (anonymous):

Let the number be written as xy, thus x*y=12 and 10x+y+36 = 10y+x solve this

OpenStudy (anonymous):

bcoz the product of the numbers is 12 hence if any one is x the other has to be 12/x

OpenStudy (anonymous):

right

OpenStudy (anonymous):

36=108/x -9x or 4=12/x -x or the reqd quad is 4x =12 -x^2 or x^2 + 4x -12 =0 or factorising (x+6)(x-2)=0 hence x=2 (neglecting -6 as it doesn't satisfy the question) hence the number is 10x+12/x =20 +6=26